OpenStudy (anonymous):

How many atoms are in 12.0 cm3 of mercury (at room temperature)?

OpenStudy (anonymous):

density of mercury at room temp is 13.6 g/cm3 as the volume is given as 12cm3, so we can calculate mass by density formula D=M/V mass=163.2 grams now calculate moles of mercury by the help of this mass MOLES=MASS/MOLar mass moles=0.816 now multiply these no of moles by avogadro's no. Ans=4.914*10^23